You are wearing a nice jersey that is 4.0 mm thick and is made of a material with thermal conductivity of 0.055 W m−1 K−1.

The inside surface of the jersey is at the same temperature as your skin, 26°C. The outside surface of the jersey is at a temperature of just 8.0°C.

If the jersey has a total surface area of 0.25 m2 then at what rate is heat being lost through the jersey? (to 2 s.f and in W)


1
Expert's answer
2021-05-18T11:11:34-0400

Rate of heat loss is given by

"Q=KA\times \frac{dT}{x}"

K = thermal conductivity of "0.055 \;W m^{−1} K^{−1}"

A = total surface area of "0.25\; m^2"

dT = difference in temperature across the thickness = 26 – 8 = 18

x = the thickness "= 4 \;mm = 4 \times ^{-3} \;m"

"Q = 0.055 \times 0.25 \times \frac{18}{4 \times 10^{-3}} = 61.875 \;W"